OpenStudy (otonogold):

Solve the system by graphing? Help?! -3x-y= -10 4x-4y=8

OpenStudy (tkhunny):

Well, graph them. Slope-Intercept Form is often convenient. Find their intersection.

OpenStudy (otonogold):

Is there a formula? @tkhunny

OpenStudy (anonymous):

y=mx+b is the formula

OpenStudy (otonogold):

how would i plug it in?

OpenStudy (anonymous):

rearrange -3x-y=-10 and 4x-4y=8 so that they are in slope intercept form: y=mx+b. do you know how to do it?

OpenStudy (otonogold):

No :/

OpenStudy (otonogold):

4y=m(4x)+8?

OpenStudy (anonymous):

\[-3x-y=-10\] \[-y=3x-10\] \[y=-3x+10\]

OpenStudy (anonymous):

how about now?

OpenStudy (anonymous):

you do the same thing with; 4x - 4y = 8 -4y = -4x - 8 ; subtract 4x from both sides y= x -2 ; divide -4 from both sides now you are left in slope intercept form y = x - 2

OpenStudy (anonymous):

are you able to graph the two? y=-3x+10 y=x-2

OpenStudy (otonogold):

hold on im graphing

OpenStudy (otonogold):

(3,-1)?

OpenStudy (anonymous):

are those the points you got from graphing?

OpenStudy (otonogold):

i dont get it

OpenStudy (mertsj):

Can you solve the two equations for y or not?

OpenStudy (otonogold):

no

OpenStudy (tkhunny):

What is your plan for graphing them? I have to believe that you are given a "solve by graphing" problem only AFTER you have been given sufficient information to graph the equations.

OpenStudy (mertsj):

Take the first equation and add 3x to both sides. What do you get?

OpenStudy (anonymous):

Y=-3x+10 all u have to do is make a table. Have x on the top of a sideways ladder tas and y on the bottom. Make 3 sections. Have -1 as x on the first section and the 0 then 1 the for y you just have to plug it in the equation and then graph it. Do the same for the other equation

OpenStudy (otonogold):

(-1,3) or (3,1)?

OpenStudy (anonymous):

Like for -1 u do -3(-1) which is 3 and thenu add10 which is 13 so your first point is (-1, 13) then do the same for the rest do u get it?

OpenStudy (otonogold):

?

OpenStudy (anonymous):

@otonoGold u have to have three x points because it doesn't give a specific coordinate so you have to pick the x coordinate except it has to be one negative number one zero then a positive of that same number like I chose -1, 0 and 1

OpenStudy (mertsj):

(3,-1) is incorrect

OpenStudy (mertsj):

So I see we are guessing now.

OpenStudy (anonymous):

Ok so I got 3 coordinates by doing what I told you to do last comment. I have (-1,13) (0,10) and (1,7) for the equation -3x-y=10

OpenStudy (anonymous):

The next equation is 4x-4y=8 like the other guy showed u we have to simplify that. So we get y= x-2 do the same table formula again an ill post my answers in a sec

OpenStudy (anonymous):

I got (-1,-3) (0,-2) and (1,-1) I would show u a pic but I don't know how to

OpenStudy (anonymous):

Hope I helped

OpenStudy (anonymous):

Then u just graph the points in a graph, but make two different graphs for the two different equations.

OpenStudy (mertsj):

Put both equations on one graph. The point where the lines cross is the solution of the system.

OpenStudy (anonymous):

Ops yeah but them in the same graph I messed up! Sorry! :( @otonogold

OpenStudy (anonymous):

Did it work?